OpenStudy (anonymous):

Artie Fact wants to determine the total value of all of the houses in his city. So he goes to the taxAssessr's Office and looks up the value of 30 houses in the city. He finds that the averages of these houses is about $87,300 houses , with a standard deviation of $5,400. From the tax rolls he finds that there are 7468 within the city limits. So he concludes that the total value of the houses is about 652 million dollars. Is this a reasonable estimate? Use the Central Value Theorem to determine your answer. Explain!

OpenStudy (anonymous):

a) Mean of the data 82+ 84 +83+ 82+ 83+ 81+ 85 +84+ 80+ 85+ 85+ 85 +87+ 83+ 82+ 83+ 84+ 83 +80 +81+ 83+ 85 +84 +80+ 79 Total= 2073 Number of items=25 =2073/25 Answer= 82.92 b) Standard Deviation of the data 82.92-82= (0.92) ^2=0.8464 82.92-80= (2.92) ^2=8.5264 82.92-84= (-1.08) ^2=1.1664 82.92-85= (-2.08) ^2=4.3264 82.92-83= (-0.08) ^2=0.0064 82.92-85= (-2.08) ^2=4.3264 82.92-82= (o.92) ^2=0.8464 82.92-85= (-2.08) ^2=4.3264 82.92-83= (-0.08) ^2=0.0064 82.92-87= (-4.08) ^2=16.6464 82.92-81= (1.92) ^2=3.6864 82.92-83= (-0.08) ^2=0.0064 82.92-85= (-2.08) ^2=4.3264 82.92-82= (0.92) ^2=0.8464 82.92-84= (-1.08) ^2=1.1664 82.92-83= (-0.08) ^2=0.0064 82.92-84= (-1.08) ^2=1.1664 82.92-83= (-0.08) ^2=0.0064 82.92-83= (-0.08) ^2=0.0064 82.92-85= (-2.08) ^2=4.3264 82.92-80= (2.92) ^2=8.5264 82.92-84= (-1.08) ^2=1.1664 82.92-81= (1.92) ^2=3.6864 82.92-80= (2.92) ^2=8.5264 82.92-79= (3.92) ^2=15.3664 Total=93.84 93.84/24=3.91 The square root of 3.91 is 1.97737

OpenStudy (anonymous):

i think this will help u a lot